OpenStudy (ohmaigosshh):

Can somebody help me? g(n - 4) = (n - 4)^2 - 5 .... would it be... n + 16 - 5 ... 16n - 5?

OpenStudy (e.mccormick):

\(g(n - 4) = (n - 4)^2 - 5\) did you try using FOIL on the \((n - 4)^2 \) part?

OpenStudy (anonymous):

First is this correct \[g(n-4)=(n-4)^{2}-5\]